The height of a triangle is 5 ft more than the length of the base. The area of the triangle is
63 ft^2 .
Find the base and height of the triangle

Answer by mananth(16946)

let base be x
height = x+5
Area = 63
1/2 * x(x+5)=63
x(x+5)=2*63
x^2+5x=126
x^2+5x-126=0
x^2+14x-9x-126=0
x(x+14)-9(x+14)=0
(x+14)(x-9)=0
x= 9ft the length of thebase
height = 9+5 =14 ft